OpenStudy (anonymous):

2205 = 315a +210c 9 = a + c

OpenStudy (anonymous):

so cost = a(315) + c(210) a is the number of adults, c is the number of children and we know there was 9 of them so a + c = 9 giving you the system 2205 = 315a +210c 9 = a + c solve one of these for a. Then replace this value of a into the other equation to find c